# InternVideo2: Scaling Video Foundation Models for Multimodal Video Understanding

### v1

@@ 5.2.4 Video Temporal Grounding
We evaluate InternVideo2 on two temporal grounding (VTG) datasets: Qvhighlight Lei et al. (2021) and Charade-STA Gao et al. (2017). The eval setting and used features are the same as in TAL. We use CG-DETR Moon et al. (2023) as the grounding head. We report R1@0.3, R1@0.5, R1@0.7, and mAP for moment retrieval as in Lei et al. (2021); Moon et al. (2023); Lin et al. (2023). Highlight Detection is evaluated in terms of “Very Good” mAP and HiT@1. In Table 13, InternVideo2-1B and InternVideo2-6B bring gradual performance improvements compared to CLIP Radford et al. (2021b) and CLIP Radford et al. (2021b)+Slowfast Feichtenhofer et al. (2019). This suggests that a larger spatiotemporal model is more beneficial to short-term video semantic alignment capabilities.
@@ 5.2.5 Audio-related Tasks
We evaluate InternVideo2’s audio and text encoders on audio tasks, including audio-text retrieval on AudioCaps Kim et al. (2019), Clothov1, and Clothov2 Drossos et al. (2020); audioQA on ClothoAQA Lipping et al. (2022) and Audio-MusicAVQA Behera et al. (2023); and audio classification on the ESC-50 Piczak (2015). As shown in Tab. 14, 15(a), and 15(b), our model achieves state-of-the-art performance on all downstream tasks. Considering the limited size of the used audio and text encoders, these audio-related results show crossmodal contrastive learning’s benefits are mutual to the used modalities. Audio and the corresponding text models also gain from this learning.
| Model | ViEncoder | LLM | MVBench | VideoChatGPT | MoVQA |
@@
| Gemini Pro Team et al. (2023) | - | - | 37.7 | - | - |
| LLaVA Liu et al. (2023) | CLIP-L | Vicuna-7B | 36.0 | - | - |
| mPLUG-Owl-V Ye et al. (2023a) | CLIP-L | LLaMA-7B | 29.7 | - | 24.8 |
| VideoChatGPT Maaz et al. (2023a) | CLIP-L | Vicuna-7B | 32.7 | 2.38 | 22.9 |
| VideoLLaMA Zhang et al. (2023c) | EVA-CLIP-g | Vicuna-7B | 34.1 | 1.98 | - |
| VideoChat2 | UMT-L | Vicuna-7B | 51.1 | 2.98 | 34.7 |
| VideoChat2 | InternVideo2-1B | Mistral-7B | 60.9 | 2.67 | 41.0 |
@@
| ViT-1B | InternVL-6B | K-Mash | 91.4 | 91.5 | 85.1 | 50.5 | 76.5 | 79.0 |
| ViT-1B | InternVL-6BVideoMAE-g | K-Mash | 91.3 | 91.4 | 85.0 | 50.8 | 77.1 | 79.1 |
| ViT-1B | InternVL-6B | K-Mash | 91.3 | 91.5 | 85.1 | 50.6 | 76.6 | 79.0 |
| ViT-6B | InternVL-6BVideoMAE-g | K-Mash | 91.9 | 91.7 | 85.7 | 51.0 | 77.5 | 79.6 |
@@ 5.3 Video-centric Dialogue and its Applications
Tab. 16 shows results on MVBench Li et al. (2023b), VideoChatGPT-Eval Maaz et al. (2023a), and MoVQA Zhang et al. (2023b) from VideoChat2 (a video-based large language model) Li et al. (2023b) with different backbones, compared with other MLLMs. Note VideoChat-InternVideo2 outperforms other systems with a notable margin, not only on the average score but also in each subtask (in the Supp), except in VideoChatGPT-Eval. Considering these benchmarks not only involve perception but also reasoning, these results suggest InternVideo2 does embed knowledge to partially model the world, at least compared with other models. It also validates the importance of learning a transferrable video representation for current video-related MLLM. We also give several qualitative evaluations with popular GPT-4V and GeminiPro, including dialogues on action sequence (Figure 4), confused action (Figure 5), temporal order understanding (Figure 6), temporal event counting (Figure 7, unexpected action reasoning (Figure 8), and vision-language navigation (Figure 9).
@@ 5.4.2 Training Data and used Teachers in Stage 1
In Tab. 18, we examine the impact of distillation teachers and used dataset sizes on model performance. (a) Data Scale: Note that pretraining data scale for MAE should grow with the increasing model scale, otherwise the downstream performance would be saturated, such as K710 (0.66M videos) for ViT-L, K-Mash for ViT-1B, and K-Mash for ViT-6B. (b) Teacher: Tab. 18 reveals that the synergy between a multimodal teacher (e.g., CLIP Radford et al. (2021b)) and a motion-aware teacher (e.g., MAE Tong et al. (2022)) markedly boosts performance, especially on SthSthV2. It highlights the importance of strategic teacher model selection in the distillation process.
@@ 6 Conclusion and Discussion
Despite its achievements, InternVideo2 does not introduce novel training methods or architectural advances. Instead, it leverages existing learning techniques for scheme exploration while focusing on improving data processing to enhance spatiotemporal perception, semantic alignment, and basic knowledge embedding. Similarly to previous studies Li et al. (2023a); Ye et al. (2023b), the model still grapples with limitations stemming from fixed input resolutions, sampling rates, and highly compressed tokens, which restrict its ability to express rich video information and capture fine-grained details.
@@ 7 Broader Impact
It is important to acknowledge that, similar to other foundational models, InternVideo2 has the potential to embed biases present in its training data and the associated models used during training, such as neural teachers Tong et al. (2022); Chen et al. (2023a) and language models (LLMs) Jiang et al. (2023); Zheng et al. (2023). These biases may emerge due to a variety of factors, including the personal ideas, preferences, values, and perspectives of the data creators and the training corpus utilized.
@@ B.1 Video-only Data
| K-Mash | 658K | 169K | 106K | 15K | 152K | 0 |
| K-Mash | 658K | 169K | 106K | 15K | 207K | 844K |
To create the curated collection of videos, named K-Mash, we source videos from renowned action recognition datasets such as Kinetics-400 (K400) Carreira and Zisserman [2017], Something-Something (Sth) Goyal et al. [2017b], Moments in Time (MIT) Monfort et al. [2020], ActivityNet Heilbron et al. [2015], and HACS Zhao et al. [2019]. These datasets provide a wide range of video types, including both first-person and third-person perspectives, short and long durations, and featuring a rich variety of characters and settings.
For the enhanced version of the dataset, called K-Mash, we push a step further and meticulously selected an additional 844,000 videos from YouTube to further enhance the diversity of the dataset. It’s important to note that all videos in this dataset are utilized for training without any labels, allowing the model to learn from the unlabeled data in an unsupervised manner. This approach helps to broaden the model’s understanding of different visual concepts and improves its performance on various video-related tasks.
@@ B.2 Videos with Audio-Video-Speech Modalities
In addtion to publicly available video-text datasets (e.g. InternVid [Wang et al., 2023d] and WebVid [Bain et al., 2021]), we introduce a new video dataset that incorporates both audio-visual-speech information and their corresponding textual descriptions. This dataset is included in the training process of InternVideo2. To create this multimodal dataset, named MVid, we leverage several video sources and provide detailed annotations. MVid includes videos with synchronized audio, visual, and speech information, along with their corresponding textual descriptions. This multimodal dataset enables the training of InternVideo2 to better understand and capture the connections between different modalities, enhancing its performance in various video-related tasks that require audio, visual, and speech understanding.
Collection. In the MVid dataset, approximately half of the videos are sourced from YouTube, while the remaining videos are gathered from anonymous sources. This is to improve the diversity of dataset, as relying solely on YouTube may result in limited depth of the dataset. Furthermore, to study the impact of video culture backgrounds on the learned models, a small portion of the dataset consists of Chinese data. These videos were collected with proper permissions for academic usage, ensuring compliance with legal and ethical considerations.
By incorporating videos from various sources and including a subset of Chinese data, MVid provides a more diverse and representative dataset for training InternVideo2. This approach allows the model to learn from a wide range of video content, encompassing different cultural backgrounds and further enhancing its ability to understand and process videos from various sources.
@@
Annotation. We automatically caption visual, audio, and speech of MVid. Then we correct them and fuse them for cross-modal captions for training using LLM. We list several annotation examples of our method in Fig. 10.
Audio Captioner. We craft a audio captioner upon VideoChat [Li et al., 2023d], as we find no reliable ones. It extracts audio features from inputs by Beats [Chen et al., 2023d]. We learn it by only tuning its Qformer (the interface between audio encoder and LLM) using a combination of the large-scale audio-text corpus WavCaps [Mei et al., 2023] dataset.
@@
Filtering & Sampling. After obtaining the captions, we calculate the CLIP similarity between the video segments and captions. We select the top 60 million data as the video segment data for MVid. For LAION-2B, we only select samples with CLIP similarity in the top 158 million for training.
B.3 Instruction-Tuning Data for Video Dialogues
We employ the training part of MVBench [Li et al., 2023b]. It comprises 1.9M samples (both images and videos) from 34 distinct sources. They are sampled from the public image / video datasets and reorganized in a standard question-anwser format. This training dataset covers key features of image and video understanding across crucial tasks, including 1) conversation, 2) caption, 3) visual question anwser, 4) reasoning, and 5) classification.
@@ C.1 Ablations
We study which part of InternVideo2’s predictions are suitable for feature-based tasks, i.e. temporal action localition. We adhere the same train and test protocols as in the main paper.
@@ C.2 Video Retrieval
We detail R@1, R@5, and R@10 of zero-shot video retrieval from InternVideo2 in Tab. 25(c) -25(f) for reference.
@@
| InternVideo2-1B | 4 | 51.9 | 74.6 | 81.7 | 49.6 | 73.6 | 81.2 |
| InternVideo2-1B | 8 | 51.9 | 75.3 | 82.5 | 50.9 | 73.4 | 81.8 |
| InternVideo2-6B | 4 | 54.5 | 77.5 | 83.7 | 52.3 | 75.3 | 83.5 |
| InternVideo2-6B | 8 | 55.9 | 78.3 | 85.1 | 53.7 | 77.5 | 84.1 |
@@
| InternVideo2-1B | 4 | 31.5 | 51.3 | 59.5 | 27.1 | 44.8 | 51.8 |
| InternVideo2-1B | 8 | 32.0 | 52.4 | 59.4 | 27.3 | 44.2 | 51.6 |
| InternVideo2-6B | 4 | 34.8 | 54.0 | 61.6 | 30.1 | 48.0 | 55.0 |
| InternVideo2-6B | 8 | 33.8 | 55.9 | 62.2 | 30.1 | 47.7 | 54.8 |
@@
| InternVideo2-1B | 4 | 70.7 | 93.7 | 96.9 | 85.9 | 97.6 | 99.2 |
| InternVideo2-1B | 8 | 70.4 | 93.4 | 96.9 | 85.4 | 97.6 | 99.1 |
| InternVideo2-6B | 4 | 71.1 | 93.8 | 97.0 | 85.2 | 97.7 | 99.4 |
| InternVideo2-6B | 8 | 71.5 | 94.0 | 97.1 | 85.3 | 97.9 | 99.3 |
@@
| InternVideo2-1B | 4 | 56.7 | 78.7 | 83.9 | 54.4 | 74.4 | 80.6 |
| InternVideo2-1B | 8 | 57.0 | 80.0 | 85.1 | 54.3 | 77.2 | 83.5 |
| InternVideo2-6B | 4 | 56.2 | 77.6 | 83.6 | 53.2 | 76.8 | 82.7 |
| InternVideo2-6B | 8 | 57.9 | 80.0 | 84.6 | 57.1 | 79.9 | 85.0 |
@@
| InternVideo2-1B | 4 | 56.9 | 81.7 | 89.8 | 53.6 | 80.0 | 88.5 |
| InternVideo2-1B | 8 | 60.4 | 83.9 | 90.8 | 54.8 | 81.5 | 89.5 |
| InternVideo2-6B | 4 | 59.4 | 83.2 | 90.3 | 53.7 | 80.5 | 88.9 |
| InternVideo2-6B | 8 | 63.2 | 85.6 | 92.5 | 56.5 | 82.8 | 90.3 |
@@
| InternVideo2-1B | 4 | 58.9 | 83.0 | 88.7 | 83.6 | 94.8 | 97.0 |
| InternVideo2-1B | 8 | 58.1 | 83.0 | 88.4 | 83.3 | 94.3 | 96.9 |
| InternVideo2-6B | 4 | 59.8 | 84.2 | 89.7 | 82.5 | 94.6 | 97.2 |
| InternVideo2-6B | 8 | 59.3 | 84.4 | 89.6 | 83.1 | 94.2 | 97.0 |

### v2

@@ 5.2.2 Video Temporal Grounding
We evaluate InternVideo2 on two video temporal grounding (VTG) datasets: Qvhighlight Lei et al. (2021), and Charade-STA Gao et al. (2017). The eval setting and used features are the same as in TAL. We use CG-DETR Moon et al. (2023) as the grounding head. We report R1@0.3, R1@0.5, R1@0.7, and mAP for moment retrieval as in Lei et al. (2021); Moon et al. (2023); Lin et al. (2023). Highlight Detection is evaluated in terms of “Very Good” mAP and HiT@1. In Table 11(b), InternVideo2-1B and InternVideo2-6B bring gradual performance improvements compared to CLIP Radford et al. (2021a) and CLIP Radford et al. (2021a)+Slowfast Feichtenhofer et al. (2019). This suggests that a larger spatiotemporal model is more beneficial to short-term video semantic alignment capabilities.
@@ 5.2.3 Audio-related Tasks
We evaluate InternVideo2’s audio and text encoders on audio tasks, including audio-text retrieval on AudioCaps Kim et al. (2019), Clothov1, and Clothov2 Drossos et al. (2020); audioQA on ClothoAQA Lipping et al. (2022) and Audio-MusicAVQA Behera et al. (2023); and audio classification on the ESC-50 Piczak (2015). As shown in Tab. 12, 13(a), and 13(b), our model achieves state-of-the-art performance on all downstream tasks. Considering the limited size of the used audio and text encoders, these audio-related results show crossmodal contrastive learning’s benefits are mutual to the used modalities. Audio and the corresponding text models also gain from this learning.
| Model | ViEncoder | LLM | MVBench | Egoschema | Perception Test |
@@
| Gemini 1.0 Pro Team et al. (2023) | - | - | 37.7 | 55.7 | 51.1 |
| Gemini 1.0 Ultra Team et al. (2023) | - | - | - | 61.5 | 54.7 |
| Gemini 1.5 Pro Team et al. (2023) | - | - | - | 72.2 | - |
| LLaVA-Next-Video Liu et al. (2024) | CLIP-L | Vicuna-7B | 46.5 | 43.9 | 48.8 |
| VideoLLaMA2 Cheng et al. (2024) | CLIP-L-336 | Mistral-7B | 54.6 | 51.7 | 51.4 |
| VideoLLaMA2 Cheng et al. (2024) | CLIP-L-336 | Mistral-8*7B | 53.9 | 53.3 | 52.2 |
| VideoChat2 | UMT-L | Vicuna-7B | 51.1 | - | - |
| VideoChat2 | InternVideo2s3-1B | Mistral-7B | 60.3 | 55.8 | 53.0 |
| VideoChat2-HD-F16 | InternVideo2s3-1B | Mistral-7B | 67.2 | 60.0 | 63.4 |
@@
| ViT-1B | InternVL-6B | K-Mash1.1M | 91.4 | 91.5 | 85.1 | 50.5 | 76.5 | 79.0 |
| ViT-1B | InternVL-6BVideoMAE-g | K-Mash1.1M | 91.3 | 91.4 | 85.0 | 50.8 | 77.1 | 79.1 |
| ViT-1B | InternVL-6B | K-Mash2M | 91.3 | 91.5 | 85.1 | 50.6 | 76.6 | 79.0 |
| ViT-6B | InternVL-6BVideoMAE-g | K-Mash2M | 91.9 | 91.7 | 85.7 | 51.0 | 77.5 | 79.6 |
@@ 5.3 Video-centric Dialogue and its Applications
Tab. 14 shows results on MVBench Li et al. (2023d), Egoschema Mangalam et al. (2023), and Perception Test Patraucean et al. (2023), by equipping VideoChat2 Li et al. (2023d) with our InternVideo2 encoder, compared with other MLLMs. Note VideoChat2HD-InternVideo2 outperforms other systems with a notable margin in Perception test and MVBench (F16 denotes the model is trained and tested with 16-frame inputs), except in Egoschema compared with proprietary commercial models (GPT4 & Gemini). Our method demonstrates superior short-term fine-grained video understanding compared to both open-source and commercial methods. Egoschema requires longer context harnessing, and we plan to further explore leveraging LLM for long multimodal sequence modeling. Considering these benchmarks not only involve perception but also reasoning, these results suggest InternVideo2 does embed knowledge to partially model the world. It also validates the importance of learning a transferrable video representation for current video-related MLLM. We also give several qualitative evaluations with popular GPT-4V and GeminiPro, including dialogues on action sequence (Figure 4), confused action (Figure 5), temporal order understanding (Figure 6), temporal event counting (Figure 7, unexpected action reasoning (Figure 8), and vision-language navigation (Figure 9).
@@ 5.4.2 Training Data and used Teachers in Stage 1
In Tab. 24, we examine the impact of distillation teachers and used dataset sizes on model performance. (a) Data Scale: Note that pretraining data scale for MAE should grow with the increasing model scale, otherwise the downstream performance would be saturated, such as K710 (0.66M videos) for ViT-L, K-Mash1.1M for ViT-1B, and K-Mash2M for ViT-6B. (b) Teacher: Tab. 24 reveals that the synergy between a multimodal teacher (e.g., CLIP Radford et al. (2021a)) and a motion-aware teacher (e.g., MAE Tong et al. (2022)) markedly boosts performance, especially on SthSthV2. It highlights the importance of strategic teacher model selection in the distillation process.
@@ 5.4.3 Training Arch, Method, and Data in Stage 2
| Model | MVBench | NextQA | Egoschema-full | Egoschema-subset |
| VideoChat2 w qformer inst | 59.9 | 79 | 52.9 | 65.8 |
| VideoChat2 w/o qformer inst | 60.4 (+0.5) | 78.6 (-0.4) | 55.8 (+2.9) | 66.4 (+0.6) |
5.4.4 Training and Evaluation in Stage 3
As shown in Table 19, incorporating questions (i.e., the ‘q’ in ‘QA’) into QFormer during Stage 3 training, which was found useful in Dai et al. (2023), actually harms the out-of-domain performance of the scaled-up VideoLLM. The NextQA training data is already included in the training corpus, and this is the only benchmark where the question-injected version performs better. Therefore, we believe that adding questions to QFormer during the instruction tuning stage of the scaled VideoChat model leads to some degree of overfitting.
@@ 6 Conclusion and Discussion
Despite its achievements, InternVideo2 does not introduce novel training methods or architectural advances. Instead, it leverages existing learning techniques for scheme exploration while focusing on improving data processing to enhance spatiotemporal perception, semantic alignment, and basic knowledge embedding. Similarly to previous studies Li et al. (2023e); Ye et al. (2023), the model still grapples with limitations stemming from fixed input resolutions, sampling rates, and highly compressed tokens, which restrict its ability to express rich video information and capture fine-grained details.
@@
Potential Biases.
We investigate the potential biases here. We focused on age, gender, and race distributions, as these are commonly recognized areas where bias can occur. We count keywords related to these categories in the used captions. Note that these synthetic captions may not fully reflect the truth of the corresponding videos, thereby creating a gap between our analysis and the actual reality. Here are the results of our analysis:
Age: The majority were about adults (86.99%), followed by children (12.87%) and barely any mentions of senior citizens (0.04%).
Gender: 62.04% pertained to men and 37.96% pertained to women.
Race: 56.19% are Asians, 23.04% are Black people, 14.55% are White people, 3.78% are Middle Eastern people, and 2.43% are Latin American people.
@@ 7 Broader Impact
It is important to acknowledge that, similar to other foundational models, InternVideo2 has the potential to embed biases present in its training data and the associated models used during training, such as neural teachers Tong et al. (2022); Chen et al. (2023c) and language models (LLMs) Jiang et al. (2023); Zheng et al. (2023). These biases may emerge due to a variety of factors, including the personal ideas, preferences, values, and perspectives of the data creators and the training corpus utilized.
@@ B.1 Video-only Data
| K-Mash1.1M | 658K | 169K | 106K | 15K | 152K | 0 |
| K-Mash2M | 658K | 169K | 106K | 15K | 207K | 844K |
To create the curated collection of videos, named K-Mash, we source videos from renowned action recognition datasets such as Kinetics-400 (K400) Carreira and Zisserman [2017], Something-Something (Sth) Goyal et al. [2017a], Moments in Time (MIT) Monfort et al. [2020], ActivityNet Heilbron et al. [2015], and HACS Zhao et al. [2019]. These datasets provide a wide range of video types, including both first-person and third-person perspectives, short and long durations, and featuring a rich variety of characters and settings.
For the enhanced version of the dataset, called K-Mash2M, we push a step further and meticulously selected an additional 844,000 videos from YouTube to further enhance the diversity of the dataset. It’s important to note that all videos in this dataset are utilized for training without any labels, allowing the model to learn from the unlabeled data in an unsupervised manner. This approach helps to broaden the model’s understanding of different visual concepts and improves its performance on various video-related tasks.
@@ B.2 Videos with Audio-Video-Speech Modalities
In addtion to publicly available video-text datasets (e.g. InternVid [Wang et al., 2023d] and WebVid [Bain et al., 2021]), we introduce a new video dataset that incorporates both audio-visual-speech information and their corresponding textual descriptions. This dataset is included in the training process of InternVideo2. To create this multimodal dataset, named InternVid2, we leverage several video sources and provide detailed annotations. InternVid2 includes videos with synchronized audio, visual, and speech information, along with their corresponding textual descriptions. This multimodal dataset enables the training of InternVideo2 to better understand and capture the connections between different modalities, enhancing its performance in various video-related tasks that require audio, visual, and speech understanding.
Collection. In the InternVid2 dataset, approximately half of the videos are sourced from YouTube, while the remaining videos are gathered from anonymous sources. This is to improve the diversity of dataset, as relying solely on YouTube may result in limited depth of the dataset. Furthermore, to study the impact of video culture backgrounds on the learned models, a small portion of the dataset consists of Chinese data. These videos were collected with proper permissions for academic usage, ensuring compliance with legal and ethical considerations.
By incorporating videos from various sources and including a subset of Chinese data, InternVid2 provides a more diverse and representative dataset for training InternVideo2. This approach allows the model to learn from a wide range of video content, encompassing different cultural backgrounds and further enhancing its ability to understand and process videos from various sources.
@@
Annotation. We automatically caption visual, audio, and speech of InternVid2. Then we correct them and fuse them for cross-modal captions for training using LLM. We list several annotation examples of our method in Fig. 10.
Audio Captioner. We craft a audio captioner upon VideoChat [Li et al., 2023c], as we find no reliable ones. It extracts audio features from inputs by Beats [Chen et al., 2023a]. We learn it by only tuning its Qformer (the interface between audio encoder and LLM) using a combination of the large-scale audio-text corpus WavCaps [Mei et al., 2023] dataset.
@@
Filtering & Sampling. After obtaining the captions, we calculate the CLIP similarity between the video segments and captions. We select the top 60 million data as the video segment data for InternVid2. For LAION-2B, we only select samples with CLIP similarity in the top 158 million for training.
| Model | Zero-shot | Finetuning | |
| Action Recognition | Video Retrieval | Action Recognition | |
| InternVideo2s2-1B | 55.5 | 55.0 | 73.2* |
| InternVideo2s2-6B | 56.9(+1.4) | 56.9(+1.9) | 73.6(+0.4)* |
| Model | Teacher | Data | K400 | K600 | K700 | MiT | SSV2 | Avg |
| ViT-L | CLIP-L | K710 | 90.3 | 90.4 | 83.2 | 48.0 | 74.7 | 77.3 |
| ViT-L | CLIP-L | K-Mash1.1M | 90.5 | 90.4 | 83.4 | 48.1 | 74.7 | 77.4 |
| ViT-L | CLIP-LMVD-L | K710 | 90.1 | 90.3 | 83.0 | 47.7 | 76.2 | 77.5 |
| ViT-L | CLIP-LMVD-LDINOv2-L | K710 | 89.8 | 90.2 | 83.0 | 47.3 | 76.4 | 77.3 |
| ViT-1B | InternVL-6B | K-Mash1.1M | 91.4 | 91.5 | 85.1 | 50.5 | 76.5 | 79.0 |
| ViT-1B | InternVL-6B | K-Mash2M | 91.3 | 91.5 | 85.1 | 50.6 | 76.6 | 79.0 |
| ViT-1B | InternVL-6BVideoMAE-g | K-Mash1.1M | 91.3 | 91.4 | 85.0 | 50.8 | 77.1 | 79.1 |
@@ C.1 Ablations
We study which part of InternVideo2s1’s predictions are suitable for feature-based tasks, i.e. temporal action localition. We adhere the same train and test protocols as in the main paper.
@@ C.2 Video Retrieval
We detail R@1, R@5, and R@10 of zero-shot video retrieval from InternVideo2 in Tab. 26(c) -26(f) for reference.
@@
| InternVideo2s2-1B | 4 | 51.9 | 74.6 | 81.7 | 49.6 | 73.6 | 81.2 |
| InternVideo2s2-1B | 8 | 51.9 | 75.3 | 82.5 | 50.9 | 73.4 | 81.8 |
| InternVideo2s2-6B | 4 | 54.5 | 77.5 | 83.7 | 52.3 | 75.3 | 83.5 |
| InternVideo2s2-6B | 8 | 55.9 | 78.3 | 85.1 | 53.7 | 77.5 | 84.1 |
@@
| InternVideo2s2-1B | 4 | 31.5 | 51.3 | 59.5 | 27.1 | 44.8 | 51.8 |
| InternVideo2s2-1B | 8 | 32.0 | 52.4 | 59.4 | 27.3 | 44.2 | 51.6 |
| InternVideo2s2-6B | 4 | 34.8 | 54.0 | 61.6 | 30.1 | 48.0 | 55.0 |
| InternVideo2s2-6B | 8 | 33.8 | 55.9 | 62.2 | 30.1 | 47.7 | 54.8 |
@@
| InternVideo2s2-1B | 4 | 70.7 | 93.7 | 96.9 | 85.9 | 97.6 | 99.2 |
| InternVideo2s2-1B | 8 | 70.4 | 93.4 | 96.9 | 85.4 | 97.6 | 99.1 |
| InternVideo2s2-6B | 4 | 71.1 | 93.8 | 97.0 | 85.2 | 97.7 | 99.4 |
| InternVideo2s2-6B | 8 | 71.5 | 94.0 | 97.1 | 85.3 | 97.9 | 99.3 |
@@
| InternVideo2s2-1B | 4 | 56.7 | 78.7 | 83.9 | 54.4 | 74.4 | 80.6 |
| InternVideo2s2-1B | 8 | 57.0 | 80.0 | 85.1 | 54.3 | 77.2 | 83.5 |
| InternVideo2s2-6B | 4 | 56.2 | 77.6 | 83.6 | 53.2 | 76.8 | 82.7 |
| InternVideo2s2-6B | 8 | 57.9 | 80.0 | 84.6 | 57.1 | 79.9 | 85.0 |
@@
| InternVideo2s2-1B | 4 | 56.9 | 81.7 | 89.8 | 53.6 | 80.0 | 88.5 |
| InternVideo2s2-1B | 8 | 60.4 | 83.9 | 90.8 | 54.8 | 81.5 | 89.5 |
| InternVideo2s2-6B | 4 | 59.4 | 83.2 | 90.3 | 53.7 | 80.5 | 88.9 |
| InternVideo2s2-6B | 8 | 63.2 | 85.6 | 92.5 | 56.5 | 82.8 | 90.3 |
@@
| InternVideo2s2-1B | 4 | 58.9 | 83.0 | 88.7 | 83.6 | 94.8 | 97.0 |
| InternVideo2s2-1B | 8 | 58.1 | 83.0 | 88.4 | 83.3 | 94.3 | 96.9 |
| InternVideo2s2-6B | 4 | 59.8 | 84.2 | 89.7 | 82.5 | 94.6 | 97.2 |
| InternVideo2s2-6B | 8 | 59.3 | 84.4 | 89.6 | 83.1 | 94.2 | 97.0 |
C.3 Multi-Choice Video Question Answering
We evaluate zero-shot multi-choice (MC) video QA using InternVideo2 in stage 2 on MSR-VTT and VATEX. Tab. 27 shows InternVideo2 consistently improves MC accuracy compared with previous SOTAs except on LSMDC, where it gets a comparable result with InternVideo.